# April 2026

### 05 Apr 2026 v1

@@ 4.3 Energy inequalities and Quantum Energy Inequalities
We shall confine our investigation to Minkowski spacetime in the representation of the Poincaré-invariant quasifree state, and we study the quantum version of condition (49), since we shall prove a theorem of crucial relevance for constructing our localization POVMs, which will later be refined in Theorem 6.2 by making the form of the quadratic form $b^{\mu\nu}_{f}$ even more precise through a suitable choice of the smearing function $f$. The results we shall present rely on important known achievements [FeSm08] (and see [San24] for recent extension). In spite of our restriction to flat spacetime and to a special reference vacuum state, the results presented in Proposition 4.10 below should be extendable to Hadamard quasifree states in curved spacetime. We shall take advantage of a special case of the main result achieved in [FeSm08], where the opposite convention to ours concerning the metric signature is adopted. We start from the observation that a normalized vector $\Psi\in{\mathfrak{S}}_{0}$ defines an algebraic state of Hadamard type [FeSm08]. Indeed, the Minkowski vacuum $\Omega$ is Hadamard, as is well known. Every normalized $\Psi\in{\mathfrak{S}}_{0}$ is such that the difference $\Lambda_{\Psi}-\Lambda_{\Omega}$ of the associated two-point functions defines a smooth integral kernel according to Proposition 4.4. Hence $\Psi$ is Hadamard as well, by definition. Take $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$, $\Psi\in{\mathfrak{S}}_{0}$ with $\Psi||=1$, and $Q=A^{\mu}(x)\partial_{x^{\mu}}+B(x)$, an (at most) first-order differential operator with smooth real coefficients. Following the procedure outlined in the proof of Theorem 3.1 of [FeSm08] – which is more generally valid in suitably shaped domains of globally hyperbolic spacetimes and for Hadamard states – taking into account the comments in Section 3 of [Few05], the result of the quoted theorem simplifies to [^9] I am grateful to C. Fewster for clarifying this point to me.
@@
    Above, $d\Sigma,d\Sigma^{\prime}$ are the ${\mathbb{R}}^{3}$ Lebesgue measure in spatial Minkowski coordinates adapted to the $3$-planes $\Sigma$ and $\Sigma^{\prime}$, respectively.
@@ 5 Relativistic spatial localization observable from the stress-energy operator
5.1 $n$-particle relativistic spatial localization from $\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}]$ satisfying CC
@@
Evident physical issues with (60) arise immediately in our QFT context. First of all, ${\mathsf{A}}_{f}^{u}(\Delta)$ defined as above is not necessarily positive, since $\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]$ is not positive, not even if we replace $f$ by $f^{2}$! However, some lower bounds hold, as shown in Proposition 4.10 when using $f^{2}$ with $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$. Furthermore, due to Proposition 4.7, if we consider the compressions to $n$-particle spaces ${{\cal H}}^{(n)}$, namely the operators given by $P_{n}:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]P_{n}$ and $P_{n}A_{f}^{u}(\Delta)P_{n}$, they are positive, and the latter is also normalized to $I$ for $\Delta=\Sigma$ because of (59) ($P_{n}{\mathfrak{F}}_{s}({{\cal H}}_{m})\to{\mathfrak{F}}_{s}({{\cal H}}_{m})$ denotes the orthogonal projector onto ${{\cal H}}^{(n)}$). We are particularly interested in the specific case of one particle: $n=1$. From the causality properties proved in Proposition 4.7 and (c) in Proposition 4.11, exploiting a procedure already used in [Mor23, DRM24, CDM26], we expect that the causality condition (CC) in Definition 2.2 is satisfied when (a compression of) the operators ${\mathsf{A}}_{f}^{u}(\Delta)$ is used to define families of POVMs on all rest spaces of $\mathbb{M}$}. Finally, notice that in (60) $\frac{1}{\sqrt{H^{u}}}\Omega$ is not defined. To fix this problem we consider the square root of the resolvent $H^{-1}_{u,\epsilon}$, where henceforth we define the $\epsilon$-regularized Hamiltonian
@@
We are now in a position to state and prove one of the main results of this work. We construct a positive-energy relativistic spatial localization observable (Definition 2.1) for states with a definite number of particles in the Fock space and which is causal in the sense that it complies with CC as in Definition 2.2. This family of POVMs uniquely arises from the stress-tensor operator smeared with a test function $f^{2}$ and the choice of a preferred temporal direction $u$. We cannot directly use an identity such as (60) to define an effect because of the states with “negative probability” that would arise from (e) in Proposition 5.1. Even if we shall return later to that issue, what we do now is remove these annoying states by passing to the compressions $P_{n}A_{f}^{u}(\Delta)P_{n}$, as already suggested.
@@
    if $\rho\in{\mathsf{S}}({{\cal H}}^{(n)})$ for a given $n=0,1,2,\ldots$ and $\Delta\in\mathscr{B}_{0}(\Sigma)$, for a spacelike $3$-plane $\Sigma$,
@@
With this definition, (a) and (b) are true, in particular because $P_{n}{\mathsf{A}}_{0,3}(\Delta)={\mathsf{A}}_{0,3}(\Delta)P_{n}$ and $P_{0}{\mathsf{A}}_{0,3}(\Delta)={\mathsf{A}}_{0,3}(\Delta)P_{0}=0$. We stress that ${A}^{u}_{f}(\Delta)\in{\mathfrak{B}}({\mathfrak{F}}_{s}({{\cal H}}_{m}))$ and that this operator is defined for every $\Delta\in\mathscr{B}(\Sigma)$ of every spacelike $3$-plane, as observed in the proof of Proposition 5.1 (see the discussion below (119) and (125)). For future convenience we note that, if $\Psi\in{{\cal H}}^{(n)}\cap{\mathfrak{S}}_{0}$ with $n>0$, we have
@@
as $\epsilon\to 0$. This is a direct consequence of the dominated convergence theorem, when one expands the scalar products as integrals on $\mathsf{V}_{m,+}^{n}$ using the fact that the function $\Psi$ is of Schwartz type (when viewed as a function of the spatial momenta $(\vec{p}_{1},\ldots,\vec{p}_{n})$), and that $(E^{n}_{\epsilon,u})^{-1/2}$ is a bounded function of $(\vec{p}_{1},\ldots,\vec{p}_{n})$. On the other hand, integration by parts proves that, for every $N=0,1,\ldots$, there are polynomials $Q_{N}(z_{1},\ldots,z_{n})$ of degree $N$ in the variables $z_{k}\in{\mathbb{R}}$, such that
@@
We conclude the proof by establishing (f) and (g). The positive-energy requirment is trivially valid due to the positivity of the Hamiltonian $H^{v}$ for $v\in{\mathsf{T}}_{+}$. Covariance of the constructed families of POVMs easily follows from (c) of Proposition 5.1 and the fact that $U$ leaves the spaces ${{\cal H}}^{(0)}$ invariant. The only non-trivial property is CC. Since ${{\cal H}}^{(0)\perp}\cap{\mathfrak{S}}_{0}$ is dense in ${{\cal H}}^{(0)\perp}$ and $A_{f}^{u}(\Delta)$ is continuous, it is sufficient to prove that the family of non-negative numbers $\langle\Psi|A_{f}^{u}(\Delta)\Psi\rangle$ satisfies CC for $\Psi\in{{\cal H}}^{(0)\perp}\cap{\mathfrak{S}}_{0}$ whose $N$ non-vanishing components are denoted by $\Psi_{n}$. We know that it holds $\langle\Psi|A_{f}^{u}\Psi\rangle=\sum_{n=1}^{N}\int_{\Delta}\langle\frac{1}{\sqrt{H^{u}}}\Psi_{n}|u^{\mu}:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]\frac{1}{\sqrt{H^{u}}}\Psi_{n}\rangle u^{\nu}_{\Sigma}d\Sigma(x)$. The proof relies entirely on the properties of the smooth current which appears in the integrand above ${\mathbb{M}}\ni x\mapsto J_{\Psi}^{\nu}(x):=\sum_{n=1}^{N}\langle\frac{1}{\sqrt{H^{u}}}\Psi_{n}|u^{\mu}:\thinspace{\hat{T}_{\mu}}\>^{\nu}\thinspace:[f^{2}_{x}]\frac{1}{\sqrt{H^{u}}}\Psi_{n}\rangle$. It is causal and future-directed wherever it does not vanish, due to Proposition 4.7, and conserved in view of (c) of Proposition 4.11. The proof of the validity of CC is then the same as for Theorems 35 and 39 in [Mor23]. ∎
@@ 5.2 Center of energy, Newton-Wigner position observable, Heisenberg inequality
A technical summary of the properties of the NW localization observable for a scalar particle and the problems it raises with causality is given in [Mor23]. We only stress that, in spite of its appealing properties, the NW localization observable cannot be considered a physically sound notion of spatial localization due to its conflict with elementary causality requirements like CC. However, as said above, it could still coincide with the first moment of several more meaningful unsharp notions of localization which satisfy CC (Proposition 61 in [DRM24]):
@@
Referring to the notions introduced in Theorem 5.2, if $\Sigma$ is any $3$-dimensional rest space, then
@@
defines a POVM on every $\Sigma$ when $\psi\in{{\cal H}}^{(1)}={{\cal H}}_{m}$, and the family of these POVMs is a positive-energy relativistic spatial localization observable for a single particle which satisfies CC.
@@ 5.3 Large-mass/non-relativistic limit: Von Neumann unsharp position measurement
    If, conversely, the function $g_{f}$ tends to become a constant function (with constant integral $1$), we obtain a more and more imprecise notion of localization in the large-mass limit. That is because the convolution $\chi_{\Delta}*g_{f}$ with an almost constant function makes indistinguishable the characteristic functions of a pair of distinct sets $\Delta$ and $\Delta^{\prime}$.
@@ 6 Commutativity of conditional localization POVMs of causally separated laboratories
(More precisely, these POVMs have the intepretation of conditional POVMs for states whose proability to find the system in $\Delta_{0}$ is close to $1$ according to (5) and the discussion below it.) As discussed in [Mor26], there is the possibility that effects of this sort, referred to different regions $\Delta_{0},\Delta^{\prime}_{0}$, included in causally separated neighborhoods, commute.
@@ 6.2 Conditional localization POVMs in laboratories
Before proving the theorem, we observe that if $\Delta_{0}$ is sufficiently regular that $L(\Delta_{0})$ is a globally hyperbolic spacetime in its own right with $\Delta_{0}$ as a smooth spacelike Cauchy surface, then the POVM $B_{\Delta_{0}}(\Delta)^{u}_{f,\eta}$ can be extended to a weaker version of relativistic spatial localization observable on the whole spacetime $L(\Delta_{0})$. This can be done with the same technology developed in [DRM24]. That mathematical technology, starting from (a) of Theorem 6.2, should also prove that the constructed relativistic spatial localization observable satisfies a natural generalization of the causal condition CC denoted by GCC therein. If $\Delta_{0}$ is only Borel, a similar generalization should be possible in the corresponding element of the causal logic according to [CDM26], proving therein the validity of an even more general causal condition for achronal sets. These issues will be investigated elsewhere. $\blacksquare$
@@ 6.4 Commutativity of conditional POVMs as local AHK operators
We remind the reader that a closed densely-defined operator $A$ is said to be affiliated with a von Neumann algebra ${\mathfrak{R}}\subset{\mathfrak{B}}({{\cal H}})$ [Tak02, StZs19] if $UA\subset AU$ for every unitary [^14] That is equivalent to the apparently stronger requirement $UAU^{\dagger}=A$ for every $\in{\mathfrak{R}}^{\prime}. UU\in{\mathfrak{R}}^{\prime}$. It turns out, by the very definition of von Neumann algebra and the fact that the unitaries of a von Neumann algebra generate the algebra, that if $A\in{\mathfrak{B}}({{\cal H}})$, affiliation with $\mathfrak{R}$} is equivalent to $A\in{\mathfrak{R}}$. It is not difficult to prove that, if $A:D(A)\to{{\cal H}}$ is selfadjoint, then $A$ is affiliated with $\mathfrak{R}$} if and only the projectors $P^{(A)}(E)$ of the PVM of $A$ are elements of $\mathfrak{R}$}.
@@
where we used Proposition 4.5. Since $\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]\Psi\in{\mathfrak{S}}_{0}$ because $\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]({\mathfrak{S}}_{0})\subset{\mathfrak{S}}_{0}$ ((a) of Proposition 4.2), the limits of both sides exist as $N\to+\infty$, and this proves the first assertion because $\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]$ – which is defined on ${\mathfrak{S}}_{0}$ – is closable. In particular, we also find that $\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}W(f)\Psi=W(f)\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\Psi$ if $\Psi\in{\mathfrak{S}}_{0}$. To prove the second assertion, take $\Phi\in D\left(\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\right)$. There is a sequence ${\mathfrak{S}}_{0}\ni\Psi_{n}\to\Phi$ such that $\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]\Psi_{n}\to\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\Phi$ by definition of closure, so that $W(f)\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\Psi_{n}\to W(f)\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\Phi$. On the other hand, $W(f)\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\Psi_{n}=\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}W(f)\Psi_{n}$, and the limit of both sides exists by construction and thus must coincide with $\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}W(f)\Phi$, since the operator under consideration is closed. In summary, if $\Phi\in D\left(\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\right)$, then $\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}W(f)\Phi=W(f)\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\Phi$. This is (b). The last statement is obvious if one observes that $\overline{A+cI}=\overline{A}+cI$ and $D(A+cI)=D(A)$, for a closable operator $A$ and any constant $c\in{\mathbb{C}}$. ∎
@@
Let us start with the case $c=0$. A densely defined closed operator $A$ is affiliated with a von Neumann algebra $\mathfrak{R}$} if $UA\subset AU$ for every unitary $U\in{\mathfrak{R}}^{\prime}$. In our case we can exploit Haag duality, ${\mathfrak{W}}({\cal O})^{\prime}={\mathfrak{W}}({\cal O}^{c})$. A unitary $U\in{\mathfrak{W}}({\cal O})^{\prime}={\mathfrak{W}}({\cal O}^{c})$ is the strong limit of finite linear combinations of Weyl operators converging to $U$, with smearing functions supported in ${\cal O}^{c}$, due to (100). From Lemma 6.15, for $\Psi\in D\left(\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]}\right)$, it holds that $W_{n}\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]}\Psi=\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]}W_{n}\Psi$. Since $\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]}$ is closed and $W_{n}\to U$ strongly, taking the limit as $n\to+\infty$ we find the desired relation $U\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]}\Psi=\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]}U\Psi$. The case $c\neq 0$ is obvious once one observes that $\overline{A+cI}=\overline{A}+cI$ for a closable operator $A$ and any constant $c\in{\mathbb{C}}$. ∎

### 07 Apr 2026 v2

@@ 4.3 Energy inequalities and Quantum Energy Inequalities
We shall confine our investigation to Minkowski spacetime in the representation of the Poincaré-invariant quasifree state, and we study the quantum version of condition (49), since we shall prove a theorem of crucial relevance for constructing our localization POVMs, which will later be refined in Theorem 6.2 by making the form of the quadratic form $b^{\mu\nu}_{f}$ even more precise through a suitable choice of the smearing function $f$. The results we shall present rely on important known achievements [FeSm08] (and see [San24] for recent extension). In spite of our restriction to flat spacetime and to a special reference vacuum state, the results presented in Proposition 4.10 below should be extendable to Hadamard quasifree states in curved spacetime. We shall take advantage of a special case of the main result achieved in [FeSm08], where the opposite convention to ours concerning the metric signature is adopted. We start from the observation that a normalized vector $\Psi\in{\mathfrak{S}}_{0}$ defines an algebraic state of Hadamard type [FeSm08]. Indeed, the Minkowski vacuum $\Omega$ is Hadamard, as is well known. Every normalized $\Psi\in{\mathfrak{S}}_{0}$ is such that the difference $\Lambda_{\Psi}-\Lambda_{\Omega}$ of the associated two-point functions defines a smooth integral kernel according to Proposition $4.$ Hence $\Psi$ is Hadamard as well, by definition. Take $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$, $\Psi\in{\mathfrak{S}}_{0}$ with $\Psi||=1$, and $Q=A^{\mu}(x)\partial_{x^{\mu}}+B(x)$, an (at most) first-order differential operator with smooth real coefficients. Following the procedure outlined in the proof of Theorem 3.1 of [FeSm08] – which is more generally valid in suitably shaped domains of globally hyperbolic spacetimes and for Hadamard states – taking into account the comments in Section 3 of [Few05], the result of the quoted theorem simplifies to [^9] I am grateful to C. Fewster for clarifying this point to me.
@@
    Above, $d\Sigma,d\Sigma^{\prime}$ are the ${\mathbb{R}}^{3}$ Lebesgue measure in spatial Minkowski coordinates adapted to the 33-planes $\Sigma$ and $\Sigma^{\prime}$, respectively.
@@ 5 Relativistic spatial localization observable from the stress-energy operator
5.1 $n$-particle relativistic spatial localization from $\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}]$ satisfying $C$
@@
Evident physical issues with (60) arise immediately in our QFT context. First of all, ${\mathsf{A}}_{f}^{u}(\Delta)$ defined as above is not necessarily positive, since $\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]$ is not positive, not even if we replace $f$ by $f^{2}$! However, some lower bounds hold, as shown in Proposition 4.10 when using $f^{2}$ with $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$. Furthermore, due to Proposition 4.7, if we consider the compressions to $n$-particle spaces ${{\cal H}}^{(n)}$, namely the operators given by $P_{n}:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]P_{n}$ and $P_{n}A_{f}^{u}(\Delta)P_{n}$, they are positive, and the latter is also normalized to II for $\Delta=\Sigma$ because of (59) ($P_{n}{\mathfrak{F}}_{s}({{\cal H}}_{m})\to{\mathfrak{F}}_{s}({{\cal H}}_{m})$ denotes the orthogonal projector onto ${{\cal H}}^{(n)}$). We are particularly interested in the specific case of one particle: $n=1$. From the causality properties proved in Proposition 4.7 and (c) in Proposition 4.11, exploiting a procedure already used in [Mor23, DRM24, CDM26], we expect that the causality condition ($C$) in Definition 2.2 is satisfied when (a compression of) the operators ${\mathsf{A}}_{f}^{u}(\Delta)$ is used to define families of POVMs on all rest spaces of $\mathbb{M}$}. Finally, notice that in (60) $\frac{1}{\sqrt{H^{u}}}\Omega$ is not defined. To fix this problem we consider the square root of the resolvent $H^{-1}_{u,\epsilon}$, where henceforth we define the $\epsilon$-regularized Hamiltonian
@@
We are now in a position to state and prove one of the main results of this work. We construct a positive-energy relativistic spatial localization observable (Definition 2.1) for states with a definite number of particles in the Fock space and which is causal in the sense that it complies with $C$ as in Definition $2.$ This family of POVMs uniquely arises from the stress-tensor operator smeared with a test function $f^{2}$ and the choice of a preferred temporal direction $u$. We cannot directly use an identity such as (60) to define an effect because of the states with “negative probability” that would arise from (e) in Proposition 5.1. Even if we shall return later to that issue, what we do now is remove these annoying states by passing to the compressions $P_{n}A_{f}^{u}(\Delta)P_{n}$, as already suggested.
@@
    if $\rho\in{\mathsf{S}}({{\cal H}}^{(n)})$ for a given $n=0,1,2,\ldots$ and $\Delta\in\mathscr{B}_{0}(\Sigma)$, for a spacelike 33-plane $\Sigma$,
@@
With this definition, (a) and (b) are true, in particular because $P_{n}{\mathsf{A}}_{0,3}(\Delta)={\mathsf{A}}_{0,3}(\Delta)P_{n}$ and $P_{0}{\mathsf{A}}_{0,3}(\Delta)={\mathsf{A}}_{0,3}(\Delta)P_{0}=0$. We stress that ${A}^{u}_{f}(\Delta)\in{\mathfrak{B}}({\mathfrak{F}}_{s}({{\cal H}}_{m}))$ and that this operator is defined for every $\Delta\in\mathscr{B}(\Sigma)$ of every spacelike 33-plane, as observed in the proof of Proposition 5.1 (see the discussion below (119) and (125)). For future convenience we note that, if $\Psi\in{{\cal H}}^{(n)}\cap{\mathfrak{S}}_{0}$ with $n>0$, we have
@@
as $\epsilon\to 0$. This is a direct consequence of the dominated convergence theorem, when one expands the scalar products as integrals on $\mathsf{V}_{m,+}^{n}$ using the fact that the function $\Psi$ is of Schwartz type (when viewed as a function of the spatial momenta $(\vec{p}_{1},\ldots,\vec{p}_{n})$), and that $(E^{n}_{\epsilon,u})^{-1/2}$ is a bounded function of $(\vec{p}_{1},\ldots,\vec{p}_{n})$. On the other hand, integration by parts proves that, for every $N=0,1,\ldots$, there are polynomials $Q_{N}(z_{1},\ldots,z_{n})$ of degree NN in the variables $z_{k}\in{\mathbb{R}}$, such that
@@
We conclude the proof by establishing (f) and (g). The positive-energy requirment is trivially valid due to the positivity of the Hamiltonian $H^{v}$ for $v\in{\mathsf{T}}_{+}$. Covariance of the constructed families of POVMs easily follows from (c) of Proposition 5.1 and the fact that $U$ leaves the spaces ${{\cal H}}^{(0)}$ invariant. The only non-trivial property is $C$. Since ${{\cal H}}^{(0)\perp}\cap{\mathfrak{S}}_{0}$ is dense in ${{\cal H}}^{(0)\perp}$ and $A_{f}^{u}(\Delta)$ is continuous, it is sufficient to prove that the family of non-negative numbers $\langle\Psi|A_{f}^{u}(\Delta)\Psi\rangle$ satisfies $C$ for $\Psi\in{{\cal H}}^{(0)\perp}\cap{\mathfrak{S}}_{0}$ whose NN non-vanishing components are denoted by $\Psi_{n}$. We know that it holds $\langle\Psi|A_{f}^{u}\Psi\rangle=\sum_{n=1}^{N}\int_{\Delta}\langle\frac{1}{\sqrt{H^{u}}}\Psi_{n}|u^{\mu}:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]\frac{1}{\sqrt{H^{u}}}\Psi_{n}\rangle u^{\nu}_{\Sigma}d\Sigma(x)$. The proof relies entirely on the properties of the smooth current which appears in the integrand above ${\mathbb{M}}\ni x\mapsto J_{\Psi}^{\nu}(x):=\sum_{n=1}^{N}\langle\frac{1}{\sqrt{H^{u}}}\Psi_{n}|u^{\mu}:\thinspace{\hat{T}_{\mu}}\>^{\nu}\thinspace:[f^{2}_{x}]\frac{1}{\sqrt{H^{u}}}\Psi_{n}\rangle$. It is causal and future-directed wherever it does not vanish, due to Proposition 4.7, and conserved in view of (c) of Proposition 4.11. The proof of the validity of $C$ is then the same as for Theorems 35 and 39 in [Mor23]. ∎
@@ 5.2 Center of energy, Newton-Wigner position observable, Heisenberg inequality
A technical summary of the properties of the NW localization observable for a scalar particle and the problems it raises with causality is given in [Mor23]. We only stress that, in spite of its appealing properties, the NW localization observable cannot be considered a physically sound notion of spatial localization due to its conflict with elementary causality requirements like $C$. However, as said above, it could still coincide with the first moment of several more meaningful unsharp notions of localization which satisfy $C$ (Proposition 61 in [DRM24]):
@@
Referring to the notions introduced in Theorem 5.2, if $\Sigma$ is any 33-dimensional rest space, then
@@
defines a POVM on every $\Sigma$ when $\psi\in{{\cal H}}^{(1)}={{\cal H}}_{m}$, and the family of these POVMs is a positive-energy relativistic spatial localization observable for a single particle which satisfies $C$.
@@ 5.3 Large-mass/non-relativistic limit: Von Neumann unsharp position measurement
    If, conversely, the function $g_{f}$ tends to become a constant function (with constant integral 11), we obtain a more and more imprecise notion of localization in the large-mass limit. That is because the convolution $\chi_{\Delta}*g_{f}$ with an almost constant function makes indistinguishable the characteristic functions of a pair of distinct sets $\Delta$ and $\Delta^{\prime}$.
@@ 6 Commutativity of conditional localization POVMs of causally separated laboratories
(More precisely, these POVMs have the intepretation of conditional POVMs for states whose proability to find the system in $\Delta_{0}$ is close to 11 according to (5) and the discussion below it.) As discussed in [Mor26], there is the possibility that effects of this sort, referred to different regions $\Delta_{0},\Delta^{\prime}_{0}$, included in causally separated neighborhoods, commute.
@@ 6.2 Conditional localization POVMs in laboratories
Before proving the theorem, we observe that if $\Delta_{0}$ is sufficiently regular that $L(\Delta_{0})$ is a globally hyperbolic spacetime in its own right with $\Delta_{0}$ as a smooth spacelike Cauchy surface, then the POVM $B_{\Delta_{0}}(\Delta)^{u}_{f,\eta}$ can be extended to a weaker version of relativistic spatial localization observable on the whole spacetime $L(\Delta_{0})$. This can be done with the same technology developed in [DRM24]. That mathematical technology, starting from (a) of Theorem 6.2, should also prove that the constructed relativistic spatial localization observable satisfies a natural generalization of the causal condition $C$ denoted by GCC therein. If $\Delta_{0}$ is only Borel, a similar generalization should be possible in the corresponding element of the causal logic according to [CDM26], proving therein the validity of an even more general causal condition for achronal sets. These issues will be investigated elsewhere. $\blacksquare$
@@ 6.4 Commutativity of conditional POVMs as local AHK operators
We remind the reader that a closed densely-defined operator AA is said to be affiliated with a von Neumann algebra ${\mathfrak{R}}\subset{\mathfrak{B}}({{\cal H}})$ [Tak02, StZs19] if $UA\subset AU$ for every unitary [^14] That is equivalent to the apparently stronger requirement $UAU^{\dagger}=A$ for every $\in{\mathfrak{R}}^{\prime}. UU\in{\mathfrak{R}}^{\prime}$. It turns out, by the very definition of von Neumann algebra and the fact that the unitaries of a von Neumann algebra generate the algebra, that if $A\in{\mathfrak{B}}({{\cal H}})$, affiliation with $\mathfrak{R}$} is equivalent to $A\in{\mathfrak{R}}$. It is not difficult to prove that, if $A:D(A)\to{{\cal H}}$ is selfadjoint, then AA is affiliated with $\mathfrak{R}$} if and only the projectors $P^{(A)}(E)$ of the PVM of AA are elements of $\mathfrak{R}$}.
@@
where we used Proposition 4.5. Since $\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]\Psi\in{\mathfrak{S}}_{0}$ because $\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]({\mathfrak{S}}_{0})\subset{\mathfrak{S}}_{0}$ ((a) of Proposition 4.2), the limits of both sides exist as $N\to+\infty$, and this proves the first assertion because $\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]$ – which is defined on ${\mathfrak{S}}_{0}$ – is closable. In particular, we also find that $\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}W(f)\Psi=W(f)\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\Psi$ if $\Psi\in{\mathfrak{S}}_{0}$. To prove the second assertion, take $\Phi\in D\left(\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\right)$. There is a sequence ${\mathfrak{S}}_{0}\ni\Psi_{n}\to\Phi$ such that $\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]\Psi_{n}\to\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\Phi$ by definition of closure, so that $W(f)\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\Psi_{n}\to W(f)\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\Phi$. On the other hand, $W(f)\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\Psi_{n}=\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}W(f)\Psi_{n}$, and the limit of both sides exists by construction and thus must coincide with $\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}W(f)\Phi$, since the operator under consideration is closed. In summary, if $\Phi\in D\left(\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\right)$, then $\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}W(f)\Phi=W(f)\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\Phi$. This is (b). The last statement is obvious if one observes that $\overline{A+cI}=\overline{A}+cI$ and $D(A+cI)=D(A)$, for a closable operator AA and any constant $c\in{\mathbb{C}}$. ∎
@@
Let us start with the case $c=0$. A densely defined closed operator AA is affiliated with a von Neumann algebra $\mathfrak{R}$} if $UA\subset AU$ for every unitary $U\in{\mathfrak{R}}^{\prime}$. In our case we can exploit Haag duality, ${\mathfrak{W}}({\cal O})^{\prime}={\mathfrak{W}}({\cal O}^{c})$. A unitary $U\in{\mathfrak{W}}({\cal O})^{\prime}={\mathfrak{W}}({\cal O}^{c})$ is the strong limit of finite linear combinations of Weyl operators converging to $U$, with smearing functions supported in ${\cal O}^{c}$, due to (100). From Lemma 6.15, for $\Psi\in D\left(\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]}\right)$, it holds that $W_{n}\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]}\Psi=\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]}W_{n}\Psi$. Since $\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]}$ is closed and $W_{n}\to U$ strongly, taking the limit as $n\to+\infty$ we find the desired relation $U\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]}\Psi=\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]}U\Psi$. The case $c\neq 0$ is obvious once one observes that $\overline{A+cI}=\overline{A}+cI$ for a closable operator AA and any constant $c\in{\mathbb{C}}$. ∎